The manager of Perrotti’s Pizza collects data concerning customer complaints. Using the information below, draw a Pareto chart to identify the critical few problems. Based on the completed Pareto chart, what is the cumulative percentage for the first two most critical issues? Problem Frequency Topping is stuck to box lid 17 Pizza arrives late 35 Wrong topping or combinations 9 Wrong style of crust 6 Wrong size 4 Pizza is partially eaten 2 Pizza never arrives 8 Q14. The cumulative percentage for the single most critical issue is ______ percent. (Enter your response as a nearest integer.) Q15. The cumulative percentage for the first two most critical issues is ______ percent. (Enter your response as a nearest
Answer #1

CAUSE

ASSIGNED

FREQUENCY

PERCENTAGE

TOPPING TRUCK IS STUCK TO BOX LID

A

17

20.99

PIZZA ARRIVES LATE

B

35

43.21

WRONG TOPPING OR COMBINATIONS

C

9

11.11

WRONG STYLE OF CRUST

D

6

7.41

WRONG SIZE

E

4

4.94

PIZZA IS PARTIALLY EATEN

F

2

2.47

PIZZA NEVER ARRIVES

G

8

9.88

TOTAL DEFECTS = 81

% OF DEFECT = INDIVIDUAL DEFECT / TOTAL DEFECT * 100

CUMULATIVE PERCENTAGE FOR MOST DEFECTS = 35 / 81 * 100 = 43.21

CUMULATIVE PERCENTAGE FOR 2 LARGEST DEFECTS = (17 + 35) / 81 * 100 = 64.19

Note: Please like the answer, if it is helpful to you. Please mention the reason in the comment section, before disliking. Thank You.......!